### v1

@@ 6.2 Comparison of Testing Power
Now, we compute and compare the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$. Specifically, we investigate how the degree of internalization, $\phi$, as well as the degree of collusion, $F/F_{c}$, influence these two statistics. Recall that for both statistics, a statistically significant positive value (greater than 1.65 at the 0.05 significance level) indicates collusion, while a significant negative value (less than -1.65 at the same significance level) indicates competition.
Table 2 reports these two statistics under various Monte Carlo configurations elaborated in Section 5.1. We first examine results without a random coefficient presented in the upper panel. Overall, our test statistic, $T_{IV}^{RV}$, performs relatively well compared to the existing statistic, $T_{markup}^{RV}$, in detecting collusion. In contrast, the latter tends to outperform the former in rejecting collusion in favor of competition at lower values of $\phi$. For example, when all six firms collude ($F_{c}=1$) and fully internalize the profits of other firms ($\phi=1$), both tests reject the competition hypothesis: $T_{IV}^{RV}=3.309$ and $T_{markup}^{RV}=3.564$. As the degree of internalization decreases (lower $\phi$), $T_{IV}^{RV}$ also declines but remains positive until $\phi=0.4$, while $T_{markup}^{RV}$ remains positive until $\phi=0.7$. When the internalization parameter is 0.1, only $T_{markup}^{RV}$ has a significantly negative value (-6.088), allowing rejection of collusion. Importantly, our test procedure performs even better in cases where $F_{c}\in\{2,3,4\}$, compared to the $F_{c}=1$ scenario, in detecting collusive behaviors among firms across a wider range of the internalization parameter $\phi$. When $F_{c}=1$, the market behaves as a single monopoly. As a result, there is no within-market variation in the IVs, and the identification power of the IVs arises solely from cross-market variations. This suggests the performance of our test may improve as the number of available markets increases.
@@
Internalization $T_{IV}^{RV}$ ($\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$) $T_{markup}^{RV}$ ($\phi=0$ vs $\phi=1$)
degree ($\phi$) $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$
@@
Notes: The table reports the median values of the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$, computed across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=6,T=100,\phi,F_{c}$). $T_{markup}^{RV}$ is constructed under the two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full internalization under the ownership structure consistent with the true effective competitors index). The top panel presents the results when a random coefficient is excluded from the indirect utility function (15) in the DGP, while the bottom panel presents the results when it is included in the utility function.
The relative testing power of our method is higher when a random coefficient is included in the consumer utility function. For any degree of collusion ($F/F_{c}$), it detects collusion effectively at high values of $\phi$ while rejecting it in favor of competition at lower values of $\phi$. However, the inclusion of a random coefficient tends to reduce the testing power of $T_{markup}^{RV}$. These results highlight the advantage of our approach: it performs comparably to, or outperforms, the existing method under various Bertrand-Nash collusive scenarios, while its implementation does not require the estimation of the demand and supply system.
Figure LABEL:fig:_p-value_iv_vuong in the Appendix presents the median p-values under the alternative hypothesis, $H_{2}:Q_{comp}>Q_{coll}$, derived from additional extensive simulations. Clearly, there is a positive monotonic relationship between testing power and the degree of internalization across all $(F,F_{c})$ configurations considered in our Monte Carlo study. The ratio $F/F_{c}$ also influences testing power; except in cases where all firms in the market collude, a higher ratio (and, therefore, a higher degree of collusion) corresponds to greater power in detecting collusive behavior among firms.
So far, we have treated the product attribute $x_{jt}$ as an exogenous variable. However, this assumption may not hold in situations where firms determine product attributes and prices simultaneously with the contemporaneous unobservable components of demand, $\xi_{jt}$. In such cases, BLP-style instruments fail to satisfy the exclusion restriction, rendering them invalid and leading to inconsistent demand estimates. To evaluate the testing power of the two statistics under this model misspecification, we generate endogenous product attributes as described in Section 5.1. The true internalization parameter $\phi$ is fixed at 1, while the direction and degree of endogeneity are parameterized by $\rho\in\{-10,-5,-1,0,1,5,10\}$ as explained in Section 5.1.
@@
In sum, we obtain results that are robust to model misspecification, which is one of the key advantages of the RV test, as highlighted by duarte2024testing. Moreover, overall, our test statistic $T_{IV}^{RV}$ demonstrates higher testing power than the existing statistic $T_{markup}^{RV}$ across all specifications, except when $F_{c}=1$, where the difference is marginal.
@@
Endogeneity $T_{IV}^{RV}$ ($\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$) $T_{markup}^{RV}$ ($\phi=0$ vs $\phi=1$)
parameter ($\rho$) $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$
@@
Notes: The table reports the median values of the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$, computed across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=6,T=100,\rho,F_{c}$). The product attribute $x_{jt}$ is treated as an endogenous variable. The direction and degree of endogeneity are parameterized by $\rho\in\{-10,-5,-1,0,1,5,10\}$, while the true internalization parameter $\phi$ is fixed at 1. $T_{markup}^{RV}$ is constructed under the two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full internalization under the ownership structure consistent with the true effective competitors index). The top panel presents the results when a random coefficient is excluded from the indirect utility function (15) in the DGP, while the bottom panel presents the results when it is included in the utility function.
Our simulation setting so far has assumed a scenario in which researchers have sufficient observations from many markets ($T=100$). To evaluate the performance of the two testing procedures when only a few markets are available in the data, we repeat our analysis with $T=10$. Results summarized in Tables LABEL:tab:_exo_vuong_few and LABEL:tab:_endo_vuong_few in the Appendix reveal that, as expected, both statistics experience a substantial loss in testing power under this data limitation. Interestingly, $T_{IV}^{RV}$ continues to yield significantly positive statistics across various collusive scenarios, whereas $T_{markup}^{RV}$ retains significant power only under a few specific configurations: positive endogeneity of $x_{jt}$ ($\rho\geq 5$), full collusion ($F_{c}=1$), and the inclusion of a random coefficient in the consumer utility, as shown in the bottom panel of Table LABEL:tab:_endo_vuong_few. The results also reveal that $T_{markup}^{RV}$ continues to outperform $T_{IV}^{RV}$ in rejecting collusion when the true industry conduct is closer to competition (low $\phi$).
@@ 6.3 Alternative functional forms for BLP-style instruments
We obtain results that largely align with our findings so far. Notably, our test statistic $T_{IV}^{RV}$, whose testing power remains relatively unaffected by changes in functional forms, continues to outperform $T_{markup}^{RV}$ in detecting collusive behavior among firms across various collusive scenarios. Interestingly, the testing power of $T_{markup}^{RV}$ declines substantially when only Local and Quadratic IVs are used but is restored when the third-order analogue is incorporated within Differentiation IVs. We also find similar effects of model misspecification arising from the endogenous product attribute $x_{jt}$ to those presented in Table 3. Additionally, the use of Differentiation IVs enhances the accuracy of the estimation for the non-linear coefficient $\sigma_{x}$, consistent with the findings of gandhi2019measuring. Further details on the alternative functional form design and extensive discussions regarding the results can be found in Appendix B.
@@ 6.4 Discussion
The most notable difference between our proposed testing procedure and the existing one is that ours does not require researchers to estimate the demand system or impose a marginal cost specification. Researchers are expected to obtain consistent and efficient demand estimates to ensure reliable and powerful testing with $T_{markup}^{IV}$. In contrast, $T_{IV}^{RV}$ does not rely on the statistical properties of demand estimates, thereby reducing potential challenges arising from the non-linear GMM framework and the inputs required for this task. Furthermore, since our testing method simply compares the model fit of two linear price regressions, it can be implemented in data-limited settings where researchers lack information on product market shares. Given these properties, our method can serve researchers and regulatory authorities as a preliminary tool for diagnosing collusive behavior among suspected firms in the market.
@@
Once the preliminary results indicate the presence of collusion, researchers can proceed with existing tests on industry conduct. The advantage of the existing statistic, $T_{markup}^{RV}$, is that researchers can test any model of conduct, provided the first-order profit-maximizing conditions can be derived from the model, and determine which model best fits the observed data based on the moment restrictions imposed on the supply-side model. On the other hand, our approach, while not assuming any particular supply model, aims to detect price collusion among firms grounded in Bertrand-Nash equilibrium conditions in differentiated goods markets. Therefore, these two approaches can complement each other, creating a more practical, efficient, and powerful framework for testing firm conduct.
@@
In addition to its preliminary role in detecting collusive behavior, our proposed testing procedure can also be used to construct stronger BLP-style instruments. For instance, once researchers observe evidence indicative of collusive behavior, they can construct $\mathbf{z}^{coll}$ based on the colluding firm index and use it in demand estimation instead of $\mathbf{z}^{comp}$. Improved performance in the first-stage demand estimation, achieved through the use of instruments that properly capture true firm conduct behavior, would ultimately enhance performance in the second-stage estimation, where (feasible) optimal instruments are employed (e.g., reynaert2014improving; conlon2020best).
7 The role of own- and other-firm IVs in the demand estimation
@@
Our proposed testing method for firm conduct utilizes $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$, which are functions of product attributes produced by the firm and its suspected colluding partners (e.g., summation of attributes or squared attributes). Now, we set aside the issue of collusion and focus on the role of the standard own- and other-firm instruments, $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$, in demand estimation. The simulation study presented in this section provides additional intuition and aids in interpreting the previous simulation results related to our proposed testing procedure.
@@
The Monte Carlo setup is as follows: We fix the number of products at $J=36$ and vary the number of firms $F\in\{1,2,\ldots,36\}$. Since we do not consider any type of collusive behavior among firms, $\phi$ is fixed at 0. We generate instruments $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$ as defined in equation (17). To investigate their roles in the demand estimation, we report the median F-statistics from the first-stage price regression and the median absolute errors of the estimated utility parameters across 500 simulations under the following instrument configurations: (i) $\mathbf{z}^{own}$ only, (ii) $\mathbf{z}^{other}$ only, and (iii) $\mathbf{z}^{both}=(\mathbf{z}^{own},\mathbf{z}^{other})$.
@@
The results for the utility specification without (with) a random coefficient, illustrated in the upper panel (bottom panel, respectively) of Figure 1, reveal several interesting points. [^21] Tables LABEL:tab:_own_vs_other_full and LABEL:tab:_own_vs_other_full_rc in the Appendix provide the full numeric results, including the median root-mean-squared errors (RMSE) of the estimated coefficients. First, own-firm instruments exhibit higher identification power than other-firm instruments in the first-stage price regression, as shown in Figures 1(a) and (c). Consequently, the median absolute error of the estimated price coefficient is lower when $\mathbf{z}^{own}$ is used as the instrument, as shown in Figures 1(b) and (d). These observations are consistent with the argument in Section 3 and previous literature (Berry95; gandhi2021empirical) that firms internalize the cross-price elasticities of their own products when setting profit-maximizing prices, resulting in greater explanatory power of own-firm product attributes in determining equilibrium prices. They also align with our simulation results presented in the previous section and echo the importance of correctly indexing firms based on the true conduct model.
@@
In the previous sections, we designed our Monte Carlo study and conducted simulations using only own-firm instruments, $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$, as they exhibit greater identification power compared to other-firm instruments in the first-stage price regression. In fact, our testing procedure can be extended to incorporate other-firm instruments constructed in the same manner as the own-firm instruments, based on the original firm indices and the suspected colluding firm indices. This is because excluding the attributes of colluding partners from other-firm instruments enhances their identification power; the price effect of own-firm product characteristics is opposite in direction to that of other-firm product characteristics, as outlined in Section 3.222222Simulation results available from the authors upon request are both quantitatively and qualitatively similar to those obtained using only the own-firm instruments. Additionally, it is worth noting that when testing whether all firms in the market collude ($F_{c}=1$), incorporating other-firm instruments into the proposed testing procedure is not valid, as there are no other-firm instruments if all firms are treated as a single entity. This feature results in a higher number of instruments in $\mathbf{z}^{comp}$ compared to $\mathbf{z}^{coll}$, making it more likely for our testing procedure to diagnose that the model of competition provides a better fit.
Figure 1: Own-firm vs. other-firm instruments
@@
Notes: The figure shows the median F-statistics from the first-stage price regression and the median absolute errors of the estimated utility parameters, $\hat{\alpha}$ and $\hat{\sigma}_{x}$, computed across 500 simulated datasets for each Monte Carlo configuration ($J=36,F,T=100$). The upper panel presents results for the utility specification without a random coefficient, while the bottom panel shows results with a random coefficient.
@@
Second, the bottom panel of the figure shows that while own-firm instruments are more effective in the first-stage price regression and more powerful for identifying the price coefficient, other-firm instruments are more effective for identifying the non-linear coefficient $\sigma_{x}$. [^23] Using $\mathbf{z}^{other}$ instead of $\mathbf{z}^{own}$ consistently results in a lower median absolute error and a lower median RMSE for the estimated non-linear coefficient, as shown in Table LABEL:tab:_own_vs_other_full_rc in the Appendix. These observations help explain the results in the bottom panel of Table 1, where using $\mathbf{z}^{coll}$ leads to a smaller median absolute error of the estimated non-linear coefficient, even when the degree of internalization is low. Specifically, the variations in product attributes of other firms unexpectedly included in $\mathbf{z}^{coll}$ may assist in identifying the non-linear coefficient.

### v2

@@ 6.2 Comparison of Testing Power
Now, we compute and compare the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$. Specifically, we investigate how the degree of profit internalization, $\phi$, as well as the degree of collusion, $F/F_{c}$, influence these two statistics. Recall that for both statistics, a statistically significant positive value (greater than 1.65 at the 0.05 significance level) indicates collusion, while a statistically significant negative value (less than -1.65 at the same significance level) indicates competition.
Table 2 reports these two statistics under various Monte Carlo configurations elaborated in Section 5.1. We first examine the results without a random coefficient presented in the upper panel. Overall, our test statistic, $T_{IV}^{RV}$, performs relatively well compared to the existing statistic, $T_{markup}^{RV}$, in detecting collusion. In contrast, the latter tends to outperform the former in rejecting collusion in favor of competition at lower values of $\phi$. For example, when all six firms collude ($F_{c}=1$) and fully internalize the profits of other firms ($\phi=1$), both tests reject the competition hypothesis: $T_{IV}^{RV}=3.309$ and $T_{markup}^{RV}=3.564$. As the degree of profit internalization decreases, $T_{IV}^{RV}$ also declines but remains statistically significant (and positive) for $\phi\geq 0.7$, while $T_{markup}^{RV}$ remains statistically significant only for $\phi\geq 0.9$. For $\phi=0.1$, only $T_{markup}^{RV}$ has a statistically significant negative value (-6.088), allowing the rejection of collusion. When $F_{c}=1$, there is no within-market variation in the collusion IVs, and their identification power relies solely on cross-market variations. While this suggests that the performance of our test may improve as the number of available markets increases, our testing procedure performs even better when $F_{c}\in\{2,3,4\}$ compared to the $F_{c}=1$ scenario in detecting collusive behaviors among firms across a broader range of $\phi$.
@@
$T_{IV}^{RV}$ ($\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$) $T_{markup}^{RV}$ ($\phi=0$ vs $\phi=1$)
$\phi$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$
@@
Notes: The table reports the median values of the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=6,T=100,\phi,F_{c}$). $T_{markup}^{RV}$ is constructed under the two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full profit internalization under industry conduct consistent with the effective firm index). The top panel presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel presents the results when it is included.
The relative testing power of our method is higher when a random coefficient is included in the consumer utility function. [^24] This higher testing power suggests that overall, the price effect of product attributes is more sensitive to industry conduct under the random coefficient specification than under the logit specification employed in our Monte Carlo study. For any degree of collusion ($F/F_{c}$), it effectively detects collusion at high values of $\phi$ while rejecting it in favor of competition at lower values of $\phi$. Conversely, the inclusion of a random coefficient tends to reduce the testing power of $T_{markup}^{RV}$. These results highlight the advantage of our approach: it performs comparably to, or outperforms, the existing method under various Bertrand-Nash collusive scenarios, while its implementation does not require estimating the demand and supply system.
Figure LABEL:fig:_p-value_iv_vuong in the Appendix presents the median p-values under the alternative hypothesis, $H_{2}:Q_{comp}>Q_{coll}$, derived from extensive additional simulations. Clearly, there is a positive monotonic relationship between testing power and the degree of profit internalization across all $(F,F_{c})$ configurations considered in our Monte Carlo study. The degree of collusion ($F/F_{c}$) also influences testing power; except in cases where all firms in the market collude, a higher degree corresponds to greater power in detecting collusive behavior among firms.
So far, we have treated the product attribute, $x_{jt}$, as an exogenous variable. However, this assumption may not hold in situations where firms determine product attributes and prices simultaneously with the contemporaneous unobservable components of demand, $\xi_{jt}$. In such cases, BLP-style instruments fail to satisfy the exclusion restriction, rendering them invalid and leading to inconsistent demand estimates. To evaluate the testing power of the two statistics under this model misspecification, we generate endogenous product attributes as described in Section 5.1. The true profit internalization parameter, $\phi$, is fixed at 1, while the direction and degree of endogeneity are parameterized by $\rho\in\{-10,-5,-1,0,1,5,10\}$ as explained in Section 5.1.
@@
In sum, our results are robust to model misspecification, which is one of the key advantages of the RV test, as highlighted by duarte2024testing. Moreover, our test statistic, $T_{IV}^{RV}$, exhibits higher testing power than the existing statistic, $T_{markup}^{RV}$, across all specifications except in the case of $F_{c}=1$, where the difference is only marginal.
@@
$T_{IV}^{RV}$ ($\mathbf{z}^{comp}$ vs $\mathbf{z}^{coll}$) $T_{markup}^{RV}$ ($\phi=0$ vs $\phi=1$)
$\rho$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$ $F_{c}=1$ $F_{c}=2$ $F_{c}=3$ $F_{c}=4$ $F_{c}=5$
@@
Notes: The table reports the median values of the two test statistics, $T_{IV}^{RV}$ and $T_{markup}^{RV}$, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F=6,T=100,\rho,F_{c}$). The product attribute $x_{jt}$ is treated as an endogenous variable. The direction and degree of endogeneity are parameterized by $\rho\in\{-10,-5,-1,0,1,5,10\}$, while the true profit internalization parameter $\phi$ is fixed at 1. $T_{markup}^{RV}$ is constructed under the two alternative firm conduct models: one with $\phi=0$ (competition) and the other with $\phi=1$ (full profit internalization under industry conduct consistent with the effective firm index). The top panel presents the results when a random coefficient is excluded from the indirect utility function (16) in the DGP, while the bottom panel presents the results when it is included.
Our simulation setting so far has assumed a scenario in which researchers have sufficient observations from many markets ($T=100$). To evaluate the performance of the two testing procedures when only a few markets are available in the data, we repeat our analysis with $T=10$. Results summarized in Tables LABEL:tab:_exo_vuong_few and LABEL:tab:_endo_vuong_few in the Appendix reveal that, as expected, both statistics experience a substantial loss in testing power under this data limitation. Interestingly, $T_{IV}^{RV}$ continues to yield significantly positive statistics across various collusive scenarios, whereas $T_{markup}^{RV}$ retains significant power only under a few specific configurations: positive endogeneity of $x_{jt}$ ($\rho\geq 5$), $F_{c}=1$, and the inclusion of a random coefficient in the consumer utility specification, as shown in the bottom panel of Table LABEL:tab:_endo_vuong_few. The results also indicate that $T_{markup}^{RV}$ tends to reject collusion more often than $T_{IV}^{RV}$ when the true industry conduct is closer to competition (low $\phi$).
@@ 6.3 Alternative functional forms for BLP-style instruments
Our results largely align with previous findings. Notably, our test statistic, $T_{IV}^{RV}$, whose testing power remains relatively unaffected by changes in functional forms, continues to outperform $T_{markup}^{RV}$ in detecting collusive behavior among firms across various collusive scenarios. Interestingly, the testing power of $T_{markup}^{RV}$ declines substantially when only Local and Quadratic IVs are used but is restored when the third-order analogue is incorporated within Differentiation IVs. We also observe similar effects of model misspecification arising from the endogenous product attribute $x_{jt}$ to those presented in Table 3. Additionally, the use of Differentiation IVs enhances the accuracy of estimating the non-linear coefficient $\sigma_{x}$, consistent with the findings of gandhi2019measuring. Further details on the alternative functional form design and extensive discussions of the results are provided in Appendix LABEL:app:_functional_forms.
@@ 6.4 Discussion
The most notable difference between our proposed testing procedure and the existing one is that ours does not require researchers to estimate the demand system or impose a marginal cost specification. Therefore, $T_{IV}^{RV}$ does not rely on the statistical properties of demand estimates, circumventing potential challenges associated with estimating the model by non-linear GMM and the inputs required for this task. [^27] Moreover, our approach is expected to be less dependent on the functional form of choice probabilities – such as nested logit, multivariate probit, and multiple-choice models – compared to existing methods that require demand estimation as a prerequisite. This is because our testing procedure relies solely on firm indexing and the nature of non-nested tests. In contrast, reliable and powerful testing with $T_{markup}^{IV}$ depends on obtaining consistent and efficient demand estimates. Furthermore, since our testing method simply compares the model fit of two linear price regressions, it can be implemented in data-limited settings where researchers lack information on product market shares. Given these properties, our method can serve as a preliminary tool for researchers and regulatory authorities to diagnose collusive behavior among suspected firms in the market.
@@
Once preliminary results indicate the presence of collusion, researchers can proceed with existing tests on industry conduct. The advantage of the existing statistic, $T_{markup}^{RV}$, is that researchers can test any model of conduct, provided the first-order profit-maximizing conditions can be derived from the model, and determine which model best fits the observed data based on the moment restrictions imposed on the supply-side model. On the other hand, our approach, while not assuming any particular supply model, aims to detect price collusion among firms grounded in Bertrand-Nash equilibrium conditions in differentiated goods markets. Therefore, these two approaches can complement each other, creating a more practical, efficient, and powerful framework for testing firm conduct.
@@
In addition to its preliminary role in detecting collusive behavior, our proposed testing procedure can also be used to construct stronger BLP-style instruments. For instance, once researchers observe evidence of collusive behavior, they can construct $\mathbf{z}^{coll}$ based on the colluding firm index and use it in demand estimation instead of $\mathbf{z}^{comp}$. Improved performance in the first-stage demand estimation, achieved through the use of instruments that properly capture true firm conduct behavior, would ultimately enhance performance in the second-stage estimation, where (feasible) optimal instruments are employed (e.g., reynaert2014improving, gandhi2019measuring, and conlon2020best).
7 The role of own- and other-firm IVs in demand estimation
@@
Our proposed testing method for firm conduct utilizes $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$, which are functions of product attributes produced by the firm and its suspected colluding partners (e.g., summation of attributes or squared attributes). Here, we set aside the issue of collusion and focus on the role of the standard own- and other-firm instruments, $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$, in demand estimation. The simulation study presented in this section provides additional intuition and aids in interpreting the previous simulation results related to our proposed testing procedure.
@@
The Monte Carlo setup is as follows: We fix the number of products at $J=36$ and vary the number of firms $F\in\{1,2,\ldots,36\}$. Since we do not consider any form of collusive behavior among firms, $\phi$ is fixed at 0. We generate instruments $\mathbf{z}^{own}$ and $\mathbf{z}^{other}$ as defined in equation (18). To investigate their roles in demand estimation, we report the median F-statistics from the first-stage price regression and the median absolute errors of the estimated utility parameters across 500 simulations under the following instrument configurations: (i) $\mathbf{z}^{own}$ only, (ii) $\mathbf{z}^{other}$ only, and (iii) $\mathbf{z}^{both}=(\mathbf{z}^{own},\mathbf{z}^{other})$.
@@
The results for the utility specification without (with) a random coefficient, illustrated in the upper panel (bottom panel, respectively) of Figure 2, reveal several interesting points. [^28] Tables LABEL:tab:_own_vs_other_full and LABEL:tab:_own_vs_other_full_rc in the Appendix provide the full numeric results, including the median root-mean-squared errors (RMSE) of the estimated coefficients. First, own-firm instruments exhibit greater identification power than other-firm instruments in the first-stage price regression, as shown in Figures 2(a) and (c). Consequently, the median absolute error of the estimated price coefficient is lower when $\mathbf{z}^{own}$ is used as the instrument, as shown in Figures 2(b) and (d). These observations are consistent with the argument in Section 3 and previous literature (Berry95; gandhi2021empirical) that firms internalize the cross-price elasticities of their own products when setting profit-maximizing prices, resulting in greater explanatory power of own-firm product attributes in determining equilibrium prices. They also align with our simulation results presented in the previous section and echo the importance of correctly indexing firms based on the true conduct model.
@@
In the previous sections, we designed our Monte Carlo study and conducted simulations using only own-firm instruments, $\mathbf{z}^{comp}$ and $\mathbf{z}^{coll}$, as they exhibit greater identification power than other-firm instruments in the first-stage price regression. In fact, our testing procedure can be extended to incorporate other-firm instruments constructed in the same manner as the own-firm instruments, based on the original firm and the suspected colluding firm indices. Excluding the product attributes of colluding partners from other-firm instruments enhances their identification power because, as outlined in Section 3, the price effects of colluding and non-colluding firms’ product attributes are opposite in direction. [^29] Simulation results available from the authors upon request are both quantitatively and qualitatively similar to those obtained using only the own-firm instruments. Additionally, it is worth noting that when testing whether all firms in the market collude ($F_{c}=1$), constructing other-firm instruments under the suspected colluding firm index is impossible, making it more likely for our testing procedure to diagnose that the model of competition provides a better fit.
Figure 2: Own-firm vs other-firm instruments
@@
Notes: The figure shows the median F-statistic from the first-stage price regression and the median absolute errors of the estimated utility parameters, $\hat{\alpha}$ and $\hat{\sigma}_{x}$, across 500 simulated datasets for each Monte Carlo configuration ($J=36,F,T=100$). The upper panel presents results for the utility specification without a random coefficient, while the bottom panel shows results with a random coefficient.
@@
Second, the bottom panel of the figure shows that while own-firm instruments are more effective in the first-stage price regression and more powerful for identifying the price coefficient, other-firm instruments are more useful for identifying the non-linear coefficient $\sigma_{x}$. [^30] Using $\mathbf{z}^{other}$ instead of $\mathbf{z}^{own}$ consistently results in a lower median absolute error and a lower median RMSE for the estimated non-linear coefficient, as shown in Table LABEL:tab:_own_vs_other_full_rc in the Appendix. These observations help explain the results in the bottom panel of Table 1, where using $\mathbf{z}^{coll}$ leads to a smaller median absolute error for the estimated non-linear coefficient, even when the degree of profit internalization is low. Specifically, variations in product attributes of other firms unintentionally included in $\mathbf{z}^{coll}$ may assist in identifying the non-linear coefficient.